OpenStudy (anonymous):

At a certain time during the day a tall tree casts a shadow of 32 feet. A shorter tree 12 feet tall casts a shadow of 16 feet. What is the height of the taller tree? 28 feet 24 feet 20 feet 16 feet

OpenStudy (lgbasallote):

use ratio and proportion the tall tree casts a shadow of 32 feet so the ratio of the height of the tree to the shadow is x feet : 32 feet the shorter tree 12 feet tall casts a shadow of 16 feet so 12 feet : 16 feet now use proportion x : 32 = 12 : 16

OpenStudy (lgbasallote):

do you get that? or do you need more help?

OpenStudy (anonymous):

more help please.

OpenStudy (anonymous):

no wait would it be 16?

OpenStudy (lgbasallote):

no

OpenStudy (lgbasallote):

so more help?

OpenStudy (anonymous):

yes i guess so lol

OpenStudy (lgbasallote):

so i can rewrite x : 32 = 12 : 16 as \[\large \frac{x}{32} = \frac{12}{16}\] because they are ratios and proportion now cross multiply \[\large 16x = 32 \times 12\] \[\large 16x = 384\] got it?

OpenStudy (anonymous):

ah yes. soo its 24?

OpenStudy (lgbasallote):

yup!

OpenStudy (lgbasallote):

nice

OpenStudy (anonymous):

so just always set it up as a potportion?

OpenStudy (lgbasallote):

well it depends on the problem..but considering as it's your lesson..then yeah

OpenStudy (anonymous):

ohkay thanks

OpenStudy (lgbasallote):

<tips hat>